OpenStudy (anonymous):

Hey! Okay, I'm working with a two circles and I need to find the standard form of another circle that is circumcentered to the first and tangent to the second. I have the slope of the line and the standard formula's of the given triengles but I don't know how to find the other point on the 2nd circle.

OpenStudy (amistre64):

your gonna have to send a pic i think

OpenStudy (anonymous):

haha! Yeah, It's kind of hard to explain. I can type in the complete problem on the sheet but I don;t have a scanner.

OpenStudy (amistre64):

use 'paint' on your computer...if you got windows that is

OpenStudy (anonymous):

ahh...okay, one moment...

OpenStudy (amistre64):

like this :)

OpenStudy (anonymous):

Okay, the circle on the bottom is the one that needs to be circumscribed and tangent to the other one on the outter side.

OpenStudy (anonymous):

I know the standard formulas for the two circles here but I need the standard fromuls for the circle that circumscribes the bottom circle and is tangent to the top

OpenStudy (amistre64):

kinda like this?

OpenStudy (anonymous):

yes!:)

OpenStudy (amistre64):

whats equations of your circles then

OpenStudy (amistre64):

you mention triangles, care to ellaborate?

OpenStudy (anonymous):

well....not exactly triengles. The tengents of a circle....the line that intersects it at only one point. Top circle is....81=(x+4)^)squared)+(y-1)^squared bottom= 64=(x+2)squared+(y+4)squared

OpenStudy (anonymous):

I'm so lost!

OpenStudy (amistre64):

lets equate these 2 and see where they intersect

OpenStudy (anonymous):

but they overlap, no?

OpenStudy (amistre64):

if there is a solution to equating them then yes...so lets check it :)

OpenStudy (amistre64):

if i did it right; they overlap at the line 4x-10y=20.... and that is providing i didnt mess it up lol

OpenStudy (anonymous):

i have the answer and it says the equation is 484=(x+2)squared+ (y+4)squared

OpenStudy (amistre64):

heres your 2 circles plotted on the graph

OpenStudy (amistre64):

and this would be your answer once the number are input

OpenStudy (amistre64):

if what you told me is right; then the new equation should equal between 196 and 225